OpenStudy (anonymous):

If y=-7 Which solution is true? A. y + 5 > –2 B. y + 5 ≥ –2 C. y + 5 < –2 D. y + 5 ≤ –2

OpenStudy (anonymous):

nevermind. it has the weird questionmark symbols

OpenStudy (phi):

refresh (reload) the page

OpenStudy (anonymous):

it's B so y>=-7

OpenStudy (anonymous):

but what would the difference between B and D be? they are both equal to, right? B is greater than or equal to, D is less than or equal to. they both have equal to!

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@felavin

OpenStudy (phi):

both B and D are true. So either you mark both, or the question has a mistake.

OpenStudy (anonymous):

would the answer matter if there was a number line with it? i'll attatch a screen shot.

OpenStudy (anonymous):

i just said y=-7 since thats where it was on the line..... i think

OpenStudy (phi):

the number line is a different question. It has an open circle at -7 which means -7 is NOT part of the solution. (A solid dot would include -7) it shows numbers bigger than -7 are part of the solution (the solid red arrow pointing to the right) so the answer is "numbers bigger than -7"

OpenStudy (anonymous):

ohhh so it is b?

OpenStudy (phi):

to write "numbers bigger than -7" you write -7 then you write either < or >. the "small end" of the < symbol points at the smaller number, and the "fat end" points at the bigger numbers. -7 < x is one way or x > -7 is another.

OpenStudy (phi):

Because you have an open circle, you use > or < if you have a solid circle, then the answer could equal -7, and you use \( \le \) or \( \ge\)

OpenStudy (anonymous):

but, im still confused. how do i solve it if y isnt -7?

OpenStudy (phi):

It looks like they are using y for the "solutions" (not x) also they added +5 to both sides.

OpenStudy (anonymous):

what does it mean by "both sides"

OpenStudy (phi):

First, which should we use y > -7 or y < -7 ?

OpenStudy (anonymous):

uhh, >?

OpenStudy (anonymous):

but what is y?

OpenStudy (phi):

in other words, how do you translate those equations into English? y > -7 reads " numbers (given the name "y") bigger than -7 Examples: -6,-5, 0, 1, etc

OpenStudy (phi):

y is a name for "numbers that are the solution"

OpenStudy (phi):

none of your choices show y > -7 but if we add +5 to both sides: y+5 > -7 + 5 and simplify the right side

OpenStudy (anonymous):

so its A?

OpenStudy (phi):

yes

OpenStudy (anonymous):

Ohhh ok. thanks for the help (im stupid in case you havent noticed :/)

OpenStudy (phi):

Look at the number line. do you see the red arrow? it is trying to show "numbers bigger than -7" does that part make sense ?

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

so then with this question, the choices would be either B or D right?

OpenStudy (anonymous):

because of the arrows

OpenStudy (phi):

now you have to learn how to write "numbers bigger than -7" using > or < you have to write -7 and then < (remember the "small end" of the < points at the smaller number, and -7 is the smallest number) so -7 < and now give a "name" for the numbers. people use x or y (why not?). Here they use y: -7 < y now you have to match that to your choices.

OpenStudy (phi):

For your second problem It is tricker.

OpenStudy (anonymous):

ruh roh...

OpenStudy (phi):

\[ - \frac{s}{2} \ge 9 \] first notice it's greater than or EQUAL, so the answer should have a solid dot (not an open circle). Unfortunately, all the choices have a solid dot.

OpenStudy (anonymous):

isnt that good?

OpenStudy (phi):

Second, you have to be careful about minus signs. one way around the trick is to add +s/2 to both sides: \[ - \frac{s}{2} +\frac{s}{2} \ge 9 +\frac{s}{2} \\ 0 \ge 9 +\frac{s}{2} \] now add -9 to both sides: \[ 0 -9 \ge 9-9 +\frac{s}{2} \\ -9 \ge \frac{s}{2} \]

OpenStudy (phi):

\[ -9 \ge \frac{s}{2} \] to get s by itself, it's ok to multiply by a (positive) number 2: \[ -9 \cdot 2\ge \frac{s}{2}\cdot 2 \\ -18 \ge s \]

OpenStudy (phi):

once you get to this stage, you should be able to pick out the correct picture. They used "s" for the name of the numbers in the solution. Can you translate this to English: \[ -18 \ge s\]

OpenStudy (anonymous):

...

OpenStudy (anonymous):

huuuuuh?

OpenStudy (anonymous):

oh wait

OpenStudy (anonymous):

B?

OpenStudy (phi):

OS is a bit flaky... but did you read the posts up above, about this problem?

OpenStudy (phi):

Did you see how to change -s/2 >= 9 to -18 >= s ? next, what does this mean : -18 >= s in English ?

OpenStudy (anonymous):

ok im back

OpenStudy (anonymous):

yes i saw how to change it

OpenStudy (anonymous):

@phi

OpenStudy (phi):

next, what does this mean : -18 >= s in English ?

OpenStudy (anonymous):

its greater or equal to s? im still confused

OpenStudy (phi):

close. -18 is bigger or equal to "all numbers we are interested in" (s are the numbers we are interested in)

OpenStudy (phi):

notice we can write it this way: s <= -18 how do you translate that to English ?

OpenStudy (anonymous):

s is less or equal to -18? I think the answer is B?

OpenStudy (phi):

Yes, s is less or equal to -18 and that means the same thing as -18 is bigger or equal to s Agree ?

OpenStudy (anonymous):

yea

OpenStudy (phi):

now think about s is less or equal to -18 Can you name a few numbers that s could be ?

OpenStudy (anonymous):

-20 -22 -24 etc? (can you help me with 1 more after this? i think its shorter)

OpenStudy (phi):

yes. almost done with this one. which choice shows an arrow that starts at -18 and "covers" your numbers -20, -22, etc. ?

OpenStudy (anonymous):

A

OpenStudy (phi):

yes. that is the answer.

OpenStudy (anonymous):

Coolio. Heres the last one.

OpenStudy (phi):

can you get t by itself ? remember, multiplying by a negative number "switches" the >= operator.

OpenStudy (anonymous):

what do you mean by "switch it"

OpenStudy (phi):

If you look back at the previous problem -s/2 >= 9 turned into s/2 <= -9 I did it by adding -s/2 to both sides, then -9 to both sides) the short way is to multiply by -1 and "switch" >= to <=

OpenStudy (anonymous):

ohh

OpenStudy (anonymous):

so now its t/5 <= -15?

OpenStudy (phi):

minus signs are tricky: -1 < 2 (for example). if we multiply by -1 on both sides we get -1*-1 < -1*2 which becomes 1 < -2 which is not true. The answer is "don't multiply by negative numbers" or "if you do, switch the operator"

OpenStudy (phi):

so now its t/5 <= -15? yes. last step: multiply both sides by 5, and simplify

OpenStudy (anonymous):

A?

OpenStudy (phi):

yes \[ 5 \cdot \frac{t}{5} \le -15 \cdot 5\\ t \le 75 \]

OpenStudy (anonymous):

Cool. Thanks for all the help (sorry i took up alot of your time)

OpenStudy (phi):

yw